The distance D that a spring is stretched by hanging object varies directly as the weight W of the object. If a 6-kg object streches a spring 19cm, find the distance the spring is streched when the weight is 19kg.
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(57307)

): The distance D that a spring is stretched by hanging object varies directly as the weight W of the object.
---
D = kW
----
Solve for "k" using "a 6-kg object streches a spring 19cm"
19 = k*6
k = 19/6
-------------------
Equation:
D = (19/6)W
Find the distance the spring is stretched when the weight is 19kg
D = (19/6)*19
D = 60.2 cm
===================
Cheers,
Stan H.
====================

Answer by lwsshak3(6492)

The distance D that a spring is stretched by hanging object varies directly as the weight W of the object. If a 6-kg object streches a spring 19cm, find the distance the spring is streched when the weight is 19kg.
**
D=(19/6)*19==≈60.17 cm